OpenStudy (anonymous):

3 greater then or equal to x-4/2 + x/3 greater than or equal to 2 x belongs to integers.

OpenStudy (anonymous):

solve the inquation and graph it on a number line

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Abhisar

OpenStudy (amistre64):

multiply by 6 to get rid of the fractions is my idea

OpenStudy (michele_laino):

we have the subsequent steps: \[3\ge \frac{ x-4 }{ 2 }+\frac{ x }{ 3 }\ge2\] \[18\ge 3(x-4)+2x \ge12\]

OpenStudy (michele_laino):

then: \[18\ge5x-12\ge12\]

OpenStudy (amistre64):

heh, i was reading it as an ending of 2x. i spose 3 to 2 is simpler to solve

OpenStudy (michele_laino):

finally:\[6\ge x \ge \frac{ 24 }{ 5 }\]

OpenStudy (anonymous):

hwo do i graph it?

OpenStudy (michele_laino):

I think in this way: